An 8 foot ladder is leaning against a building. The angle the ladder makes with the wall is 11 degrees. How far up the building does the ladder reach? Show your work.
Answer by ikleyn(52781)

.

    How far up the building does the ladder reach = 8*cos(11°) = 8*0.9816 = 7.85 ft.    ANSWER

Solved, with the work shown.